OpenStudy (anonymous):

Please Help Me! Find tan2θ if θ terminates in Quadrant IV and cosθ = 4/5?

Nnesha (nnesha):

 first of all make a triangle according to \[\rm sin = \frac{ opposite }{ hypotenuse }~~~~ \cos = \frac{ adjacent }{ hypotenuse } ~~\tan= \frac{ opposite }{ adjacent }\] cos = adjacent /hyp now use Pythagorean theorem and find 3rd side of right triangle